Simplify: x/50 + x/25 + 3x/50 + 2x/25 + ... + 2x =

Answer by Fombitz(32388)

Factor out (x/50),
(1/50)x+(2/50)x+(3/50)x+...+(100/50)x=(x/50)(1+2+3+...+100)
(1/50)x+(2/50)x+(3/50)x+...+(100/50)x=(x/50)(5050)
(1/50)x+(2/50)x+(3/50)x+...+(100/50)x=101x